OpenStudy (anonymous):

@apple_pi simplest form of fibonacci polynomial \[\frac{1}{1-x-x^2}=F_0+F_1x+F_2x^2+...+F_kx^k+...\] \(F_k\) is the \(k\text{th}\) term of Recursive Fibonacci Sequence this series will be converge if \(|x+x^2|<1\)

OpenStudy (anonymous):

let suppose \(F_0\) , \(F_1\) , \(F_2\) ,..... are the fibonacci terms....

OpenStudy (anonymous):

we want to find a function \( f(x)\) such that :\[f(x)=F_0+F_1x+F_2x^2+....+F_kx^k+...\]

OpenStudy (experimentx):

i know the generating function!! but don't know how it works.

OpenStudy (anonymous):

\[\ \ f(x)=F_0+F_1x+F_2x^2+....\]\[xf(x)= \ \ \ \ \ \ \ F_0x+F_1x^2+F_2x^3+....\]\[x^2f(x)= \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ F_0x^2+F_1x^3+F_2x^4+....\] subtract \(xf(x)\) and \(x^2 f(x)\) from \(f(x)\) u will get \[f(x)-xf(x)-x^2f(x)\\ \ \ \ \ \ =F_0+(F_1-F_0)x+(F_2-F_1-F_0)x^2+...+(F_k-F_{k-1}-F_{k-2})x^k+...\]

OpenStudy (anonymous):

@apple_pi can u tell me \(RHS\) of last equation equals to what ... see the coefficients of polynomial

OpenStudy (anonymous):

wait, does it equal F0

OpenStudy (anonymous):

F1-F0 = 1-1 = 0 F2-F1-F0 = 2-1-1 = 0

OpenStudy (anonymous):

Yep thats right

OpenStudy (anonymous):

and because \[F_k=F_{k-1}+F_{k-2}\]

OpenStudy (anonymous):

Oh, so 1 = f(x)-xf(x)-x^2f(x)

OpenStudy (anonymous):

then divide by f(x) to get 1/f(x) = 1-x-x^2 f(x) = 1/(1-x-x^2)!

OpenStudy (anonymous):

OK thank you very much

OpenStudy (anonymous):

Right